OpenStudy (anonymous):

Solve for X: Log 2 ^16 = X

OpenStudy (anonymous):

log 2 16 = log 2 2^4= 4 log 2 2 = 4 * 1= 4 see log a b = b * log a

OpenStudy (anonymous):

and log a a=loga/loga=1

OpenStudy (anonymous):

please use link for basics of log:: http://www.purplemath.com/modules/logrules.htm

OpenStudy (raden):

you mean, like this ?  @DanaDiane1234